# TimeFlow: Longitudinal Brain Image Registration and Aging Progression Analysis

### v2

@@ 3.2 Interpolation Constraints
rigid 10.3$\pm$1.9 16.9$\pm$1.5 - - 7.3$\pm$1.1 20.6$\pm$1.5 - -
NODER 9.3$\pm$1.9 18.0$\pm$1.7 1.8$\pm$0.9 0.2$\pm$1.9 5.5$\pm$0.9 23.2$\pm$1.5 1.8$\pm$0.5 0.0$\pm$0.3
VXM 9.1$\pm$1.9 18.3$\pm$1.9 6.3$\pm$4.8 $10^{4}\pm 10^{4}$ 5.6$\pm$1.1 22.9$\pm$1.7 1.5$\pm$0.2 0.2$\pm$0.4
VXM-diff 9.2$\pm$2.0 18.4$\pm$1.9 6.3$\pm$2.0 $10^{3}\pm 10^{3}$ 5.5$\pm$1.0 23.2$\pm$1.7 2.2$\pm$0.3 0.1$\pm$0.2
TimeFlow 8.6$\pm$1.6 18.8$\pm$1.7 5.0$\pm$1.0 $10^{3}\pm 10^{3}$ 5.5$\pm$1.0 23.0$\pm$1.7 1.6$\pm$0.1 0.1$\pm$0.4
TimeFlow-diff 8.8$\pm$1.8 18.7$\pm$1.9 6.0$\pm$1.0 6.1$\pm$11.2 5.3$\pm$1.0 23.3$\pm$1.8 2.2$\pm$0.2 0.0$\pm$0.0
@@
Interpolation Consistency To encourage plausible time-continuous deformations, we propose the interpolation consistency as a regularizer:
which combines the two constraints shown in Fig. 2(b). We formulated the loss with the notation $t_{k}$, but as described in the previous paragraph, $t_{k}$ can also be sampled from the uniform distribution if no intermediate observation exists. It is derived from the common transformation composition consistency that two consecutive transforms are equivalent to a direct transform. Moving from $t=0$ to $t=1$ and then back to $t=t_{k}$ is equivalent to moving directly from $t=0$ to $t=t_{k}$ (arrow 1 composed with arrow 3 should equal to arrow 2 in Fig. 2(b)). The backward case is derived similarily. It is worth noting that this differs from the commonly used smoothness regularizations, which explicitly enforce a smooth gradient over the flow field. Therefore, there is only a minimal trade-off between accuracy and smoothness since both the similarity and the flow consistency encourage plausible and reasonable longitudinal transformations.
3.3 Extrapolation Consistency
Through the intermediate similarity and interpolation consistency, we obtain meaningful predictions for bidirectional registrations in $t\in[-1,1]$. As our main objective is extrapolation, we formulate criteria for the intervals $[1,+\infty)$ (forward) and $(-\infty,-1]$ (backward) next. To this end, we consider the deformation $\phi_{1}^{0\rightarrow k}$ from $I_{0}$ to $I_{k}$. Given the time of $I_{k}$ being $t_{k}$, it holds that $\phi_{t_{k}}^{0\rightarrow L}=\phi_{1}^{0\rightarrow k}$. As illustratedin Fig. 2(b), they both register $I_{0}$ to $I_{k}$. Consequently, we can then formulate $\phi_{1}^{0\rightarrow L}=\phi_{1/t_{k}}^{0\rightarrow k}$ (arrow 1 equals to arrow 5 in Fig. 2(b)), where $\phi_{1/t_{k}}^{0\rightarrow k}$ denotes an extrapolation because $t_{k}\in(0,1)$ and therefore $11/t_{k}>$ 1. Based on this notation, we define the extrapolation similarity as:
Again, we can either use intermediate observations with time $t_{k}$ or sample from $\mathcal{U}(0,1)$. If $t_{k}$ is sampled from $\mathcal{U}(0,1)$, i.e., $I_{k}$ is not observed but only $I_{0}$ and $I_{L}$, we then define $I_{k}=I_{0}\circ\phi_{t_{k}}^{0\rightarrow L}$. In addition to the extrapolation similarity measure, we define the forward extrapolation flow consistency as:
As shown in the bottom green part of Fig. 2(b), both $\phi_{1}^{0\rightarrow L}$ and $\phi_{1/t_{k}}^{0\rightarrow k}$ map from $I_{0}$ to $I_{L}$, thus should be consistent with each other. The backward versions of the extrapolation similarity and extrapolation flow consistency are:
The similarity measures and flow consistencies both contribute to achieving a realistic and coherent extrapolation. The proposed losses effectively ensure accurate extrapolation while implicitly capturing temporal consistency and continuity, without the need for additional temporal smoothness constraints.
4 Data and Experiments
Figure 3: Visualization of Interpolation and Extrapolation Results. TimeFlow/diff takes the Baseline and Year 4 images as input, interpolates to Year 1.5, and extrapolates to Year 8. The first row displays the ground truth images. The second row shows the moved baseline image. The third and fourth rows display error maps: the initial error map before registration and the error map after TimeFlow/diff registration. Error values range from $[-1,1]$ due to the image intensity normalization to $[0,1]$. These results demonstrates that TimeFlow/diff achieves high accuracy in both temporal interpolation and extrapolation. Detailed temporal deformations can be referred to Appendix.
Data and Pre-processing.
We evaluate our method on the public Alzheimer’s Disease Neuroimaging Initiative (ADNI) dataset, which comprises longitudinal T1-weighted Magnetic Resonance (MR) scans from both healthy subjects and patients diagnosed with Alzheimer’s Disease (AD) or MCI, i.e., neurodegenerative disorders that exhibit specific patterns of atrophy development in the brain. We employ the longitudinal processing from FreeSurfer v7.2 [13]. First, all scans are processed individually, including skull stripping, normalization. Subsequently, the image sequence is rigidly aligned via a groupwise rigid registration [39]. FreeSurfer aggregates information across time points and refines the pre-processed outputs. We first select 134 subjects who underwent substantial anatomical change (median change magnitude within the brain is $>0.3$mm) over time. Then, the dataset is divided into training, validation, and testing sets, ensuring a balance in diagnosis, age, sex, and the number of visits. It results in 74, 30, 30 subjects for each split, respectively, with the number of time points per subject ranging from 2 to 12. All input scans have isotropic 1mm spacing and are cropped to a uniform size of $160\times 160\times 192$ voxels. The intensity values are scaled to the range [0,1], from percentile [0,99.9] without clipping.
Implementation Details.
Our model is implemented using Pytorch 2.1.2. and training/test is carried out on one NVIDIA A100 40G with a batch size of $1$. We set the dimension of time embedding to $16$, the feature channels in the U-Net encoder and decoder to $[32,32,48,48,96]$. In addition to TimeFlow, which outputs the time-dependent deformation field directly, we have the diffeomorphic variant TimeFlowdiff, which utilizes the scaling and squaring layer with $7$ integration steps to integrate the temporal stationary velocity field. The weights of both interpolation and extrapolation similarity are set to $1.0$. The weights of interpolation and extrapolation flow consistency are set to $2,0.03$ for TimeFlow and $1.25,0.025$ for TimeFlowdiff.
Baselines.
To evaluate the registration performance we compare TimeFlow with the following methods:
  • NODER [51, 3], which is aware of every time point and takes the complete image sequence as the model input. It carries out subject-specific optimization and models the time-varying velocity field with a neural network and leverages Neural ordinary differential equation (ODE) [8] to compute integration and adjoint gradients. It fits the subject-specific temporal deformation trajectory by maximizing the similarity with each frame.
  • VoxelMorph (VXM) [4] and the extended diffeomorphic VoxelMorph (VXMdiff[10]. Since both are pairwise registration methods, the interpolation and extrapolation prediction is carried out based on the linear scaling of the displacement or velocity field from the pairwise registration. The integration steps of VXMdiff is also set to 7.
  • Sequence-based VoxelMorph (Seq-VXM) [50]. Seq-VXM processes the whole sequence and registers the baseline image to each intermediate time point. Given that Seq-VXM performs multiple pairwise registrations by taking the whole image sequence as model input, it is aware of all images at every time point. Therefore, we treat Seq-VXM as the performance upper bound for the retrospective study (refer to Tab. 1).
Figure 4: Visualization of the Extrapolation Results. Left: Error maps show differences between the extrapolated predictions and the ground truth future images. Right: Future image predictions for different $t$ values, where no ground truth is available. $t=1$ represents one-step pairwise registration between $I_{m}$ and $I_{f}$. Thus, $t=2,4,6$ represents two, four, six-step extrapolation over $I_{f}$. Major errors are highlighted with red arrows. VXM/diff produces unsmooth deformations, leading to structural inconsistencies and loss of anatomical details (zoom in for more details).
For all learning-based methods, we apply Local Normalized Cross Correlation (LNCC) as the training loss, the same as in TimeFlow/diff. The radius of LNCC in NODER is set to $r=9$ while for the other methods, it is set to $r=4$. The loss function weighting factors in NODER follows their original suggestion. Seq-VXM, VXM and VXMdiff compute the SOTA smoothness regularizer GradICON [45, 46] with weight $1.5,0.5,1.0$.
Evaluation metrics.
@@ 4 Data and Experiments
Figure 5: Left: Progression Analysis for Control (CN), MCI and Dementia. A progression rate closer to 1 indicates slower pathological brain structural change. Right: Relative Progression of 9 representative subjects. Each colored dot represents one visit. The $x$-axis denotes the years since the first visit. The $y$-axis stands for the relative progression rate compared to the baseline visit. A higher value indicates faster brain structural changes. The grey line in the background is the mean fitted linear line across subjects. Colormap: CN-green, MCI-orange, Dementia-red.
@@ 5.1 Prospective studies: Predicting the future
We first evaluate TimeFlow’s performance in prospective studies, i.e., for extrapolating the deformations to predict follow-up images. Our model can perform this prospective analysis in two ways: 1. Based on two available scans and given a future timepoint $t$ ($t>1$), our model can directly predict the extrapolated flow field and the warped image at $t$. 2. Given a future scan, our model can estimate the corresponding $\phi_{t}$, thus the time $t$, which can reflect the subjects’ biological aging progression.
Given t, predict the future image.
As discussed in Sec. 3, one of TimeFlow’s biggest benefits is its ability to naturally extrapolate the deformation and predict a future image given $t>1$. The quantitative evaluation of extrapolation is showcased in table Tab. 1. TimeFlow demonstrates superior extrapolation accuracy compared to all baselines with the lowest MAE and the highest PSNR. Seq-VXM is not compared here since it cannot carry out future predictions. Fig. 3 (Year 8) and Fig. 4 (Left) illustrate TimeFlow’s extrapolation prediction when ground truth images are available, while Fig. 4 (Right) demonstrates the extrapolation given a large $t$ where ground truth is not available anymore. As shown in Fig. 4, NODER extrapolates unrealistic deformations, leading to significant error in future image predictions. VXM/diff, which applies linear scaling to flow fields, results in unsmooth deformation and dissolves the brain, causing anatomical details to fade. In contrast, TimeFlow/diff generates smooth and accurate extrapolated flow fields, aligning with the lowest MAE and low NDV in Tab. 1.
Given future scan, assess the brain aging.
We can utilize the extrapolation capability of TimeFlow for disease progression analysis. For each future scan, we predict a corresponding $t$ value based on prior visits. We compare the relative brain aging rate across diagnosis groups in the ADNI dataset, i.e., Controls, MCI, and Dementia. Specifically, MRI triplets are sampled from the validation and test set, where the intervals between the first and second timepoints are similar to the intervals between the second and third timepoints. The first two time points are used as input to extrapolate to the third, future time point. By comparing with the ground truth MRI, we compute the optimal extrapolated time $t_{\textrm{ext}}$. From this, we can compute the brain aging rate by dividing the extrapolated time by the actual time interval. The idea is that if brain aging is accelerated, i.e., in MCI or Dementia patients, the progression rate will be larger than 1, meaning that brain aging is faster than chronological aging. As shown in the boxplot of Fig. 5, Dementia exhibits the highest progression rate, followed by MCI, while the control group remains is close to $1$, highlighting a biologically plausible separation of groups based with our brain aging rate metric. To visualize the brain aging rate over time, we plot subject-specific progression curves for 9 representative subjects (3 from each group). We observe accelerated progression compareed to the baseline visit in MCI and Dementia, while the control group (CN) maintains a flatter curve. This highlights how TimeFlow’s extrapolation capabilities enables meaningful prospective analysis on Alzheimer’s disease, without the need for additional anatomical annotations.
@@ 5.2 Retrospective studies: Matching the deformation trajectory
We conduct studies on timepoints interpolation to assess whether TimeFlow learns temporally continuous and plausible deformation field connecting intermediate visits. In this setting, the baseline visit $I_{0}$ and the last scan $I_{L}$ serve as inputs, and the deformation fields $\phi_{t_{k}}$ are estimated to register intermediate images $I_{k}$, where $0<K<L$. Quantitative comparisons with baseline methods are presented in Tab. 1. As discussed in Sec. 4, Seq-VXM process the entire sequence and performs explicit pairwise registration for each timepoint, making it the performance upper bound. Despite using only the baseline and last scan as inputs, TimeFlow outperforms NODER, which is also aware of the entire sequence and parametrized by NeuralODE. Additionally, TimeFlowdiff achieves the least non-diffeomorphic volume without any extra spatial smoothness regularization. The interpolation results shown in Fig. 3 (Year 1.5) futher demonstrate that TimeFlow achieves almost zero residual errors compared to the initial misalignment.
@@ 5.3 Ablation Study
As described in Sec. 3 and supported by both quantitative and qualitative results, TimeFlow achieves temporally continuous and spatially smooth deformation fields without requiring explicit temporal or spatial smoothness regularization. To assess the effectiveness of our proposed consistency losses, we conduct ablation studies. To evaluate the temporal smoothness and continuity of TimeFlow, we compute the derivative $\partial\phi_{t}/\partial t$ using the finite difference method by sampling $t$ at intervals of $\Delta t=0.25$. Similar to the commonly used diffusion regularizer in learning-based registration methods, which computes the Dirichlet energy over spatial gradients, we quantify temporal smoothness as $\|\partial\phi_{t}/\partial t\|$. Specifically, we compute the norm of the temporal displacement derivatives within the brain and then take the mean over all brain voxels as the final metric. As shown in Tab. 2, incorporating interpolation consistency loss significantly improves both temporal and spatial smoothness ($\|\partial\phi_{t}/\partial t\|$ and NDV drastically decrease) while maintaining similar interpolation accuracy. Furthermore, the extrapolation similarity and flow consistency losses enable TimeFlow to predict plausible extrapolated deformation fields for future image estimation. These results demonstrate that TimeFlow inherently learns temporally continuous and spatially smooth deformation fields without explicit regularization, thus leading to the least trade-off between accuracy and smoothness. Visualization of the temporal deformations can be referred to Appendix.
6 Discussion and Conclusion
Novelty.
@@ 6 Discussion and Conclusion
TimeFlow introduces novel interpolation and extrapolation constraints that ensure temporal consistency without temporal smoothness constraints. Together with the time-conditioning mechanism inspired by diffusion models, this enables effective extrapolation beyond the observed time frame. While we have used UNet as the base model, our contributions are generic and can, therefore, also be integrated into other registration backbones. Importantly, TimeFlow has significant application potential in areas such as in-silico drug effectiveness analysis. By extrapolating brain images to future timepoints at the beginning of a treatment, the predicted outcomes can be compared with actual follow-up scans. Discrepancies between predicted and observed results may then be attributed to treatment effects, offering a novel and efficient approach for evaluating therapeutic efficacy in clinical settings.
Limitation and failure cases.
@@
TimeFlow encounters challenges in the following two scenarios: 1. Future image prediction/aging progression analysis with minimal biological aging differences: When the two base input images are temporally close and exhibit no noticeable aging differences, TimeFlow struggles to predict future images effectively based on them. In such cases, the two input images collapse into a single timepoint, rendering the model incapable of extrapolation based on a “single” input data. 2. Prediction for large time intervals: Although TimeFlow is designed to estimate future images for any given time parameter $t\in[0,\infty]$, its performance deteriorates for large $t$ values e.g. $t$ > 6. This limitation is expected due to the insufficient availability of long-range temporal data in the ADNI dataset, meaning TimeFlow rarely encountered such scenarios during training, and, of course, the fact that modeling such long developments, which are influenced by a plethora of parameters, is inherently very challenging.
Conclusion.

### v3

@@ 4 Data and Experiments
[…]
Figure 6: Visualization of interpolation (Year 1.5) and extrapolation (Year 8) results from coronal and axial views of a subject’s brain. TimeFlow/diff takes the Baseline and Year 4 images as input, interpolates to Year 1.5, and extrapolates to Year 8. Ground truth images (row 1) and warped images (row 2) are compared. Initial misalignment (row 3) is effectively corrected (row 4), demonstrating TimeFlow’s accuracy in temporal modeling. Error values range from [1,1][-1,1] due to the image intensity normalization to $[0,1]$.
Baselines
To evaluate the registration performance, we compare TimeFlow with the following methods:
  • NODER [17, 16] takes the complete image sequence as the model input and is thus aware of every time point. It carries out subject-specific optimization and models the time-varying velocity field with a neural network and leverages NeuralODE [52] to compute integration and adjoint gradients. It fits the subject-specific temporal deformation trajectory by maximizing the similarity with each frame.
  • VoxelMorph (VXM) [25] and the diffeomorphic variant (VXMdiff[53]. Since both are pairwise registration methods, the interpolation and extrapolation prediction is carried out based on the linear scaling of the displacement or velocity field from the pairwise registration. The number of integration steps of VXMdiff is also set to 7.
  • Sequence-based VoxelMorph (Seq-VXM) [15]. Seq-VXM processes the whole sequence and registers the baseline image to each intermediate time point. Given that Seq-VXM performs multiple pairwise registrations by taking the full sequence as model input, we treat Seq-VXM as the performance upper bound for the retrospective study (refer to Tab. 1).
All methods use Local Normalized Cross Correlation (LNCC) with radius $r=9$ (NODER) or $r=4$ (others) as image similarity loss. The loss function weighting factors in NODER follow their original suggestion. Seq-VXM, VXM and VXMdiff compute the SOTA smoothness regularizer GradICON [54, 55] with weight $1.5,0.5,1.0$.
Figure 7: Left: Prospective, annotation-free analysis of brain aging progression rates for Control (CN), MCI and Dementia. Rates near 1 reflect normal, chronological brain aging, whereas values above 1 indicate accelerated neurodegenerative progression. Right: Retrospective group-wise analysis of longitudinal brain progression. Trajectories represent median linear fits between TimeFlow’s inferred temporal brain state ($t$) and chronological years since baseline. Shaded ribbons indicate the 40th-60th percentile range. A steeper slope corresponds to faster neuroanatomical changes. Colormap: CN-green, MCI-orange, Dementia-red.
Figure 8: Quantitative comparison of extrapolation accuracy among TimeFlow and baseline methods on ADNI dataset. Barplots represent mean values with standard errors. Significance levels by Wilcoxon $t$-test (Holm–Bonferroni correction) are indicated. ***:p<0.001, **:p<0.01, *:p<0.05
Evaluation metrics
[…]
@@ 5.1 Prospective studies: Predicting the future
We first evaluate TimeFlow’s performance in prospective scenarios, specifically, its ability to extrapolate deformation fields to predict future brain scans. TimeFlow supports two forms of prospective analysis: (1) Given two available scans and a future time variable $t>1$, the model directly predicts the corresponding deformation fields and resulting future MRI images; (2) Given an observed future scan, TimeFlow can infer the corresponding $\phi_{t}$ and use the estimated time variable $t$ as a surrogate metric to quantify the biological aging progression.
Given t, predict the future image
A significant advantage of TimeFlow, as described in Section˜3, is its ability to extrapolate deformation fields to future timepoints beyond the observed interval. The quantitative performance of TimeFlow’s extrapolation forecasting is reported in Table˜1. The metrics are computed with respect to the ground truth future scans. Results clearly demonstrate that TimeFlow consistently achieves superior extrapolation accuracy compared to baseline methods, yielding the lowest MAE and highest PSNR (Figure˜8), even when evaluated in a zero-shot manner on the unseen OASIS dataset. Note that Seq-VXM cannot carry out future predictions; thus, it is excluded from this comparison. Qualitative extrapolation results are visualized in Figure˜6 (Year 8) and Figure˜4, illustrating TimeFlow’s accuracy by comparing to the ground truth images. Moreover, Figure˜5 demonstrates extrapolation at a large $t$ ($t=2,4,6$) where ground truth is not available. In these cases, NODER produces unrealistic and anatomically implausible deformations, resulting in significant errors in future image prediction. Similarly, VXM/diff, which applies linear scaling to displacement/velocity fields, generates unsmooth deformations (indicated by high NDV) and dissolves the brain, causing the anatomical details to fade. In contrast, TimeFlow/diff consistently generates smooth and coherent extrapolated deformation fields, quantitatively supported by the lowest MAE and minimal NDV values in Table˜1.
Given future scan, assess the brain aging
As emphasized in Section˜1, the extrapolation capability of TimeFlow enables annotation-free disease progression analysis. We compare the brain aging rates across diagnostic groups, i.e., Controls (CN), MCI, and Dementia, by predicting the optimal time parameter ($t_{\textrm{ext}}$) corresponding to the observed future MRI scans. Specifically, MRI triplets are sampled from validation and test subsets, where the intervals between the first and second visits match closely the intervals between the second and third visits. The diagnosis at the third visit labels each triplet. Using the first two scans as input, TimeFlow extrapolates forward to the third, future scan. By comparing predictions against actual observed future, we compute the extrapolated time parameter $t_{\textrm{ext}}$, and divide it by the normalized chronological intervals (in years) to yield a brain aging rate. An aging rate above 1 indicates accelerated pathological changes, as expected in patients with neurodegenerative conditions. Indeed, the boxplot of Figure˜7 (Left) clearly shows that Dementia patients exhibit the highest progression rate, followed by MCI subjects, while the control group maintains progression rates around 1. The analysis validates that our annotation-free surrogate brain aging rate metric is able to separate disease groups.
[…]
@@ 5.2 Retrospective studies: Matching the deformation trajectory
We further evaluate TimeFlow’s interpolation capability, assessing its effectiveness in modeling temporally continuous and anatomically plausible deformation trajectories connecting observed MRI visits. In this setting, TimeFlow uses only the baseline visit $I_{0}$ and the final visit $I_{L}$ as inputs, predicting deformation fields $\phi_{t_{k}}$ for intermediate images $I_{k}$, where $0<k<L$. Quantitative comparisons with baseline methods for interpolation are presented in Table˜1. As discussed in baselines, Seq-VXM, performing explicit pairwise registrations for every available timepoint, represents a performance upper bound. Despite using only the baseline and last scan as inputs, TimeFlow outperforms NODER on the ADNI dataset, which is also aware of the entire sequence and parametrized by Neuralordinary differential equation (ODE). TimeFlow achieves comparable performance to NODER in zero-shot evaluation on the OASIS dataset, where NODER performs subject-specific optimization on the full image sequence. Additionally, TimeFlowdiff achieves the least non-diffeomorphic volume without extra spatial smoothness regularization. Qualitative visualizations in Figure˜6 (Year 1.5) illustrate the accuracy of TimeFlow’s interpolations, substantially reducing alignment errors compared to initial misalignments. Similar to our above prospective analysis, we also perform retrospective, group-wise brain progression analysis using the inferred temporal parameter $t$ of intermediate visits as a surrogate brain progression state marker. We extract trajectory segments with the same diagnoses from each subject, and then compute a linear relationship between the estimated brain state ($t$) and chronological age (years since baseline of the segment). The median linear lines of the groups (CN, MCI, Dementia) in Figure˜7 (Right) clearly illustrate distinct progression patterns, with Dementia exhibiting the fastest structural changes, MCI intermediate, and CN the slowest.
@@ 5.3 Ablation Study
As described in Section˜3 and supported by both quantitative and qualitative results, TimeFlow achieves temporally continuous and spatially smooth deformation fields without requiring explicit smoothness regularization. To assess the effectiveness of our proposed inter-/extra-polation consistency constraints, we perform detailed ablation studies. To evaluate the temporal smoothness and continuity of TimeFlow, we compute the derivative $\partial\phi_{t}/\partial t$ using the finite difference method by sampling $t$ at intervals of $\Delta t=0.25$. Similar to the commonly used diffusion regularizer in learning-based registration methods, which computes the Dirichlet energy over spatial gradients, we quantify temporal smoothness as $\|\partial\phi_{t}/\partial t\|$. Specifically, we compute the norm of the temporal displacement derivatives within the brain regions and then take the mean over all brain voxels as the final metric. As shown in table˜2, incorporating interpolation consistency loss significantly improves both temporal and spatial smoothness ($\|\partial\phi_{t}/\partial t\|$ and NDV drastically reduce) without sacrificing interpolation accuracy. Furthermore, the extrapolation similarity and flow consistency losses enable TimeFlow to generate plausible extrapolated deformation fields for future image estimation. These ablation results demonstrate that TimeFlow inherently learns temporally continuous and spatially smooth deformation fields, achieving an optimal balance between registration accuracy and smoothness without explicit smoothness penalties.
6 Discussion
Contributions
@@ 6 Discussion
TimeFlow introduces novel interpolation and extrapolation constraints that ensure temporal consistency without explicit temporal smoothness regularization. Combined with the time-conditioning mechanism, TimeFlow enables effective extrapolation beyond the observed time frame. While we have used U-Net as the base model, our contributions are generic and can be integrated into other registration backbones. Importantly, TimeFlow has promising application potential in areas such as in silico drug effectiveness analysis. By extrapolating brain images to future timepoints at the beginning of a treatment, the predicted outcomes can be compared with actual follow-up scans. Discrepancies between predicted and observed results could then be attributed to treatment effects, offering a novel and efficient approach for evaluating therapeutic efficacy in clinical settings.
Bidirectional extrapolation
While we mainly focus on the forward extrapolation (predicting future brain states) in this study, TimeFlow is inherently bidirectional and allows backward extrapolation as well. Backward extrapolation, i.e., traveling back in time and visualizing the brain at earlier, unobserved ages, opens several avenues: (i) reconstructing a pre-symptomatic baseline for patients before neurodegeneration; (ii) generating counterfactual controls that share a subject’s anatomy but simulate a healthy aging trajectory.
Limitation and failure cases
@@
TimeFlow encounters challenges in the following two scenarios: 1. Future image prediction/aging progression analysis with minimal biological aging differences: When the two input scans are temporally close and exhibit no noticeable structural differences, they collapse into a single timepoint, which TimeFlow struggles to predict future images effectively. 2. Prediction for far future: Although TimeFlow is designed to estimate future images for any given time parameter $t\in[0,\infty]$, its performance deteriorates for large $t$ values, e.g., $t$ > 6. This limitation is expected due to limited data covering such long-range time intervals in the ADNI dataset, meaning TimeFlow rarely encountered such scenarios during training. The modeling of such long developments, which are influenced by a plethora of parameters, is inherently challenging.
7 Conclusion